# A citation index bridging Hirsch’s $h$ and Egghe’s $g$

### v1

@@ Abstract
We propose a citation index $\nu$ (“nu”) and show that it lies between the classical $h$-index and $g$-index. This idea is then generalized to a monotone parametric family $(\nu_{\alpha})$ ($\alpha\geq 0$), whereby $h=\nu_{0}$ and $\nu=\nu_{1}$, while the limiting value $\nu_{\infty}$ is expressed in terms of the maximum citation.
1. Introduction
Hirsch (2005) made a breakthrough in scientometrics by proposing for the first time a simple citation index (commonly referred to as $h$-index), which had an advantage of aggregating the author’s productivity on the basis of both the number of published papers and their quality. Before that, only some extensive summary statistics were used, such as the mean number of citations per paper. Since then, the $h$-index has become a standard metric of authors’ reputation and productivity, for instance routinely taken into consideration in academic appointments and promotions.
@@ 1. Introduction
To remedy such censoring of larger citations, the alternative citation index (referred to as $g$-index) was proposed by Egghe (2006a), defined as the maximum number $g$ of an author’s most cited papers, such that their total number of citations is at least $g^{2}$. From this definition, it is easy to see that $h\leq g$ (Egghe, 2006b).
These two (by now classical) indexes have attracted a lot of interest and generated ample research into their analytic properties and performance on real datasets, including their estimation in a variety of statistical models of count data (see, e.g., Egghe & Rousseau (2006); Egghe (2006b); Egghe & Rousseau (2008); Yong (2014)). Furthermore, many modifications and alternative variants of the $h$ and $g$ indexes have been proposed, focusing on certain features of the citation profile (see, e.g., Alonso et al. (2009); Guns & Rousseau (2009); Hirsch (2019) and further references therein). An up-to-date online resource SCI2S (2025) is especially useful by keeping track of the state-of-the-art in this domain.
In the present work, we introduce a new citation index $\nu$ (“nu”) aiming to bridge the mathematical definitions of Hirsh’s $h$ and Egghi’s $g$. This idea was first coined in Nuermaimaiti (2023). Namely, we start by observing that the $h$-index can be represented as a sum of certain indicator functions that censor papers to ensure a required minimum of citations. Building on this observation, our $\nu$-index essentially mimics the summative nature of the $h$-index but the new summation explicitly involves the citation numbers of the top papers.
@@
$h\leq\nu\leq g^{*},$ (1)
where $g^{*}$ denotes a modified (unconstrained) $g$-index, obtained if we are allowed to add fictitious zeros to the citation vector (Woeginger, 2008b).
On the other hand, a “tempered” version $\bar{\nu}$ of the $\nu$-index, modified so as to be not larger than the number of published papers, satisfies the inequalities
$h\leq\bar{\nu}\leq g.$ (2)
We will finish off by introducing a more general family of citation indexes $(\nu_{\alpha})$, where $\nu_{\alpha}$ is a non-decreasing (integer-valued) function of a real parameter $\alpha\geq 0$. Here, $h=\nu_{0}$ and $\nu=\nu_{1}$, while the limiting value $\nu_{\infty}$ can be expressed in terms of the maximum citation. Concluding remarks will touch upon evaluation of citation indexes in specific frequency models using the concept of limit shape (Nuermaimaiti et al., 2021; Bogachev et al., 2023; Nuermaimaiti, 2023; Yong, 2014).
2. The $h$ and $g$ indexes
@@
We say that a vector $\bm{x}=(x_{1},\dots,x_{m})$ is dominated by a vector $\bm{y}=(y_{1},\dots,y_{\ell})$ (written as $\bm{x}\preccurlyeq\bm{y}$) if $x_{i}\leq y_{i}$ for all $i\geq 1$; more precisely, if $m\leq\ell$ then $x_{i}\leq y_{i}$ for $i=1,\dots,m$, but if $m>\ell$ then $x_{i}\leq y_{i}$ for $i=1,\dots,\ell$ and $x_{i}=0$ for $i=\ell+1,\dots,m$. Effectively, these two cases imply that we complement the absent components of either $\bm{x}$ or $\bm{y}$ with fictitious zeros to equalize their dimensions, and then the dominance holds component-wise.
The following natural conditions are commonly assumed to be satisfied by any reasonable citation index $c(\bm{x})$ (see Woeginger (2008b); cf. Woeginger (2008a)):
@@
    If $\bm{x}=(x_{1},x_{2},\dots,x_{m})$ and $\bm{y}=(x_{1},x_{2},\dots,x_{m},0)$ then $c(\bm{x})=c(\bm{y})$.
@@
Let us now recall the above verbal definitions of the $h$ and $g$ indexes and put them into an explicit mathematical formulation.
Starting with the $h$-index, its definition can be expressed as follows,
$h\equiv h(\bm{x})=\max\left\{j\geq 1\colon\!\sum_{i=1}^{m}{\bf 1}_{\{x_{i}\geq j\}}\geq j\right\}\!,$ (3)
where the indicator notation ${\bf 1}_{A_{j}}$ is defined as ${\bf 1}_{A_{j}}=1$ if the condition $A_{j}$ is satisfied, and ${\bf 1}_{A_{j}}=0$ otherwise.
Note that the maximum in (3) is uniquely defined, since the sum on the left-hand side is a decreasing function of $j$, while the right-hand is strictly increasing.
Thus, the sum in (3), for a given $j$, represents the number of papers with at least $j$ citations, while the inequality requires that this number is not smaller than $j$. In particular, noting that the sum of indicators in (3) is not larger than $m$, it follows that the $h$-index is bounded by the number of papers:
$h\leq m.$ (4)
In the degenerate case with $x_{i}\equiv 0$ (i.e., $\bm{x}=\bm{0}$), the inequality in (3) is only satisfied for the value $j=0$, which is excluded from the testing range; thus, the resulting set of suitable $j$’s is empty and, according to the common convention, its maximum is set to be zero: $\max\varnothing=0$; hence
$h(\bm{0})=0,$ (5)
so that property (C1) is automatically satisfied. It is also easy to see that (C2) holds (because $h$ is insensitive to zero citations) and that (C3) is also true.
Next, the definition of the $g$-index can be written mathematically as follows,
$g\equiv g(\bm{x})=\max\left\{1\leq k\leq m\colon\sum_{i=1}^{k}x_{i}\geq k^{2}\right\}\!.$ (6)
@@
Denote $s_{k}=x_{1}+\dots+x_{k}$. According to definition (6), the index $g$ is the largest value of $k$ for which $s_{k}\geq k^{2}$. But it may be unclear whether the inequality $s_{k}\geq k^{2}$ can fail for some $k<g$. Let us show that $s_{k}\geq k^{2}$ for all $k\leq g$.
@@
If $s_{k}<k^{2}$ for some $k\geq 1$, then $s_{\ell}<\ell^{2}$ for all $\ell\geq k$.
@@
that is, $s_{k+1}<(k+1)^{2}$. The general claim then follows by induction. ∎
@@
If $s_{k}<k^{2}$ then $g<k$. In particular, if $s_{m}<m^{2}$ then $g<m$; otherwise, $g=m$.
@@
Turning to the verification of the required properties (C1)–(C3) for the $g$-index, we see that (C1) automatically holds due to (7). It is also easy to see that (C3) holds as well. However, the result of Lemma 2 suggests that, a bit surprisingly, property (C2) may fail. For instance, for $\bm{x}=(4)$, we have $g(\bm{x})=1$, but for $\bm{y}=(4,0,0,0)$ (i.e., with three fictitious zeros added), definition (6) yields $g(\bm{y})=2$.
To salvage (C2), and also to amplify the role of top-cited papers, it was suggested (Egghe, 2006b; Woeginger, 2008b) to lift the constraint $k\leq m$ in the definition of the $g$-index (see (6)) by complementing the citation vector $\bm{x}=(x_{1},\dots,x_{m})$ with additional zeros, as if such papers have been published. We denote this version of $g$ by $g^{*}$:
$g^{*}\equiv g^{*}(\bm{x})=\max\left\{k\geq 1\colon\sum_{i=1}^{k}x_{i}\geq k^{2}\right\}\!.$ (10)
This makes no difference if $g\leq m$, because then $g^{*}=g$. However, the case where $g=m$ may be drastically different.
@@
Suppose that $s_{m}\geq(m+1)^{2}$. Then $g^{*}=\bigl{\lfloor}\sqrt{s_{m}}\bigr{\rfloor}\geq m+1>m=g$.
@@
Adding $g^{*}-m\geq 1$ fictitious zeros to the citation vector $\bm{x}$, we get a new vector $\bm{x}^{\prime}=(x_{1},\dots,x_{m},0,\dots,0)$ of dimension $m+(g^{*}-m)=g^{*}$, resulting in
$\sum_{i=1}^{g^{*}}x^{\prime}_{i}=s_{m}\geq(g^{*})^{2},$
and it follows that $g^{*}=\bigl{\lfloor}\sqrt{s_{m}}\bigr{\rfloor}$. ∎
A striking example illustrating this situation is the case of John Nash (Woeginger, 2008b), with the (rounded) citation vector $\bm{x}=(2000,2000,1500,1000,400,250,100,100)$, for which we get $g=8$ but $g^{*}=85$.
3. An alternative citation index $\nu$
Trying to reconcile the definition of the $h$-index given by formula (3), with the definition of the $g$-index in (6) by taking into account the actual citations of the top papers, we propose a new citation index called the $\nu$-index, defined as the maximum integer $\nu$ such that the total sum of citation counts of papers with at least $\nu$ citations each is not less than $\nu^{2}$. Mathematically, this is expressed as (cf. (3))
$\nu\equiv\nu(\bm{x})=\max\left\{j\geq 1\colon\sum_{i=1}^{m}x_{i}\mbox{$\;\!$}{\bf 1}_{\{x_{i}\geq j\}}\geq j^{2}\right\}\!.$ (11)
Similarly to (3), the maximum is uniquely defined, noting that the sum in (11) is a decreasing function of $j$, while the right-hand side is strictly increasing. It is also worth pointing out that, unlike $g$ vs. $g^{*}$, the $\nu$-index is insensitive to fictitious zeros.
Simple examples show that the value of $\nu$ may be larger than the total number of papers, in contrast with the $h$ and $g$ indexes. For instance, for $\bm{x}=(9,6,1)$ we get $\nu(\bm{x})=4>3$.
@@ 3. An alternative citation index $\nu$
$\bar{\nu}\equiv\bar{\nu}(\bm{x})=\max\left\{1\leq j\leq m\colon\sum_{i=1}^{m}x_{i}\mbox{$\;\!$}{\bf 1}_{\{x_{i}\geq j\}}\geq j^{2}\right\}\!.$ (12)

### v2

@@ Abstract
We propose a new citation index $\nu$ (“nu”) and show that it lies between the classical $h$-index and $g$-index. This idea is then generalized to a monotone parametric family $(\nu_{\alpha})$ ($\alpha\geq 0$), whereby $h=\nu_{0}$ and $\nu=\nu_{1}$, while the limiting value $\nu_{\infty}$ is expressed in terms of the maximum citation.
Keywords: citation indexes, $h$-index, $g$-index, scientometrics, citation data
Significance Statement
The widely used Hirsch’s $h$-index values productivity but overlooks how highly each paper is cited, while Egghe’s $g$-index emphasizes top-cited work but neglects lower-cited contributions. To address these imbalances, we propose the $\nu$-index, a synthetic metric that accounts for both highly and modestly cited publications and, therefore, offers a fairer and more balanced assessment of research impact.
1 Introduction
1.1 Background
Hirsch (2005) made a breakthrough in scientometrics by proposing for the first time a simple citation index (commonly referred to as $h$-index), which had the advantage of aggregating the author’s productivity on the basis of both the number of published papers and their quality measured by generated citations. Before that, only some extensive summary statistics were used, such as the mean number of citations per paper. Since then, the $h$-index has become a standard metric of authors’ reputation and productivity, for instance routinely taken into consideration in academic appointments and promotions.
@@ 1.1 Background
To remedy such censoring of larger citations, an alternative citation index (referred to as $g$-index) was proposed by Egghe (2006a), defined as the maximum number $g$ of an author’s most cited papers, such that their total number of citations is at least $g^{2}\mbox{$\;\!\!$}$. From this definition, it is easy to see that $h\leq g$ (Egghe, 2006b).
These two (by now classical) indexes have attracted a lot of interest and generated ample research into their analytic properties and performance on real datasets, including their estimation in a variety of statistical models of count data (see, e.g., Egghe, 2006b; Egghe & Rousseau, 2006, 2008; SCI2S, 2025). Furthermore, many modifications and alternative variants of the $h$ and $g$ indexes have been proposed, focusing on certain features of the citation profile (see, e.g., Alonso et al., 2009; Guns & Rousseau, 2009; Hirsch, 2019; SCI2S, 2025, and further references therein).
1.2 New index and layout
In the present work, we introduce a new citation index $\nu$ (“nu”) aiming to bridge the mathematical definitions of Hirsch’s $h$ and Egghe’s $g$. This idea was first coined in Nuermaimaiti (2023). Namely, we start by observing that the $h$-index can be represented as a sum of certain indicator functions that censor papers to ensure a required minimum of citations. Building on this observation, our $\nu$-index essentially mimics the summative nature of the $h$-index but the new summation explicitly involves the numbers of citations of the top papers.
@@ 1.2 New index and layout
$h\leq\nu\leq g^{*}\!,$
where $g^{*}\!$ denotes a modified (unconstrained) $g$-index, obtained if we are allowed to add fictitious zeros to the citation vector (Woeginger, 2008b). On the other hand, a “tempered” version $\bar{\nu}$ of the $\nu$-index, modified so as to be not larger than the number of published papers, satisfies the inequalities
$h\leq\bar{\nu}\leq g.$
We will finish off by introducing a more general family of citation indexes $(\nu_{\alpha})$, where $\nu_{\alpha}$ is a non-decreasing (integer-valued) function of a real parameter $\alpha\geq 0$. Here, $h=\nu_{0}$ and $\nu=\nu_{1}$, while the limiting value $\nu_{\infty}$ can be expressed in terms of the maximum citation.
1.3 Disclaimer
There have been a lot of discussions about the utility and limitations of the indexes $h$ and $g$ (see, e.g., Alonso et al., 2009; Costas & Bordons, 2007; Hirsch, 2007; SCI2S, 2025; Thelwall, 2025; Waltman, 2016, and references therein), including their questionable predictive power. In this work, we only interpret citation indexes as a suitable characteristic of productivity. However, some further thoughts about the societal dimension of citation indexes will be added in the Conclusion section.
2 The $h$ and $g$ indexes
2.1 Notation
Let us fix some notation. Suppose that an author has published $m\geq 1$ papers, with the ordered numbers of citations $x_{1}\geq\dots\geq x_{m}$, where $x_{i}\geq 0$ are integers (possibly zero). We call $\bm{x}=(x_{1},\dots,x_{m})$ the citation vector. A zero vector $\bm{0}=(0,\dots,0)$ represents the degenerate case of no citations (note that the variable dimension of this vector is determined by the number of published papers). Denote
$S_{k}=x_{1}+\dots+x_{k},\qquad k=1,\dots,m.$ (1)
Clearly, $n=S_{m}$ is the total number of citations generated by the $m$ papers. Furthermore, we write
$m_{*\mbox{$\;\!\!$}}(j)=\sum_{i=1}^{m}\bm{1}_{\{x_{i}\geq j\}}=\#\{x_{i}\geq j\}$ (2)
@@ 2.1 Notation
Following Woeginger (2008b), we say that a vector $\bm{x}=(x_{1},\dots,x_{m})$ is dominated by a vector $\bm{y}=(y_{1},\dots,y_{\ell})$ (written as $\bm{x}\preccurlyeq\bm{y}$) if $x_{i}\leq y_{i}$ for all $i\geq 1$; more precisely, if $m\leq\ell$ then $x_{i}\leq y_{i}$ for $i\leq m$, but if $m>\ell$ then $x_{i}\leq y_{i}$ for $i\leq\ell$ and $x_{i}=0$ for $\ell<i\leq m$. Effectively, these two cases imply that we complement the absent components of either $\bm{x}$ or $\bm{y}$ with fictitious zeros to equalize their dimensions, and then the dominance holds component-wise.
Remark 1.
The component-wise dominance $\bm{x}\preccurlyeq\bm{y}$ should not be confused with (weak) majorization $\bm{x}\prec_{w}\bm{y}$, defined by the conditions $\sum_{i=1}^{k}x_{i}\leq\sum_{i=1}^{k}y_{i}$, for all $k$ (see Marshall et al., 2011, pp. 11–12). Like before, the lengths of vectors $\bm{x}$ and $\bm{y}$ are equalized by adding fictitious zeros as necessary. Clearly, if $\bm{x}\preccurlyeq\bm{y}$ then $\bm{x}\prec_{w}\bm{y}$, but not conversely.
2.2 Generic properties
The following natural conditions are commonly assumed for any reasonable citation index $c(\bm{x})$ (Woeginger, 2008b) (cf. Woeginger, 2008a):
@@ 2.2 Generic properties
    If $\bm{x}=(x_{1},\dots,x_{m})$ and $\bm{y}=(x_{1},\dots,x_{m},0)$ then $c(\bm{x})=c(\bm{y})$.
@@
Remark 2.
The majorization relation $\prec_{w}$ looks more flexible as a comparative tool. The corresponding version of property (C3) is stated similarly:
  • (C3)
    If $\bm{x}\prec_{w}\bm{y}$ then $c(\bm{x})\leq c(\bm{y})$.
However (perhaps, surprisingly), the $h$-index does not satisfy (C3): e.g., for $\bm{x}=(2,2)$ and $\bm{y}=(8,1)$ we have $\bm{x}\prec_{w}\bm{y}$ but $h(\bm{x})=2>h(\bm{y})=1$. On the other hand, the $g^{*}\!$-index (but not $g$) does satisfy (C3) (see definitions (6) and (8) below); e.g., $g^{*}\mbox{$\>\!\!$}(\bm{x})=2<g^{*}\mbox{$\>\!\!$}(\bm{y})=3$.
2.3 Mathematical expressions and relations for $h$ and $g$
Let us now recall the above verbal definitions of the $h$ and $g$ indexes and put them into an explicit mathematical formulation. Starting with the $h$-index, its definition can be expressed as follows,
$h\equiv h(\bm{x})=\max\left\{j\geq 1\colon\!\sum_{i=1}^{m}\bm{1}_{\{x_{i}\geq j\}}\geq j\right\}\!,$ (3)
or, using notation (2),
$h=\max\bigl\{j\geq 1\colon m_{*\mbox{$\;\!\!$}}(j)\geq j\bigr\}.$ (4)
Note that the maximum in (3) is uniquely defined, since the sum on the left-hand side is a decreasing function of $j$, while the right-hand side of the testing inequality is strictly increasing.
In particular, noting that $m_{*\mbox{$\;\!\!$}}(j)\leq m$, it follows that the $h$-index is bounded by the number of papers:
$h\leq m.$
In the degenerate case with $x_{i}\equiv 0$ (i.e., $\bm{x}=\bm{0}$), the inequality in (3) is only satisfied for the value $j=0$, which is excluded from the testing range; thus, the resulting set of suitable $j$’s is empty and, according to the common convention, its maximum is set to be zero: $\max\varnothing=0$; hence $h(\bm{0})=0$, so that property (C1) is automatically satisfied. It is also easy to see that (C2) holds (because $h$ is insensitive to zero citations) and that (C3) is also true.
Next, the definition of the $g$-index can be written as follows,
$g\equiv g(\bm{x})=\max\left\{1\leq k\leq m\colon\sum_{i=1}^{k}x_{i}\geq k^{2}\right\}\!,$ (5)
or, recalling notation (1),
$g=\max\left\{1\leq k\leq m\colon S_{k}\geq k^{2}\right\}\!.$ (6)
@@ 2.3 Mathematical expressions and relations for $h$ and $g$
2.4 Auxiliary lemmas for sums
According to definition (6), the index $g$ is the largest value of $k\leq m$ for which $S_{k}\geq k^{2}$. But it may be unclear whether the inequality $S_{k}\geq k^{2}$ can fail for some $k<g$. Let us show that $S_{k}\geq k^{2}$ for all $k\leq g$.
@@ 2.4 Auxiliary lemmas for sums
If $S_{k}<k^{2}\mbox{$\>\!\!$}$ for some $k\geq 1$, then $S_{\ell}<\ell^{2}\mbox{$\>\!\!$}$ for all $\ell\geq k$.
@@
that is, $S_{k+1}<(k+1)^{2}\mbox{$\;\!\!$}$. The general claim then follows by induction. ∎
@@
If $S_{k}<k^{2}\mbox{$\>\!\!$}$ then $g<k$. In particular, $g<m$ or $g=m$ according as $S_{m}<m^{2}\mbox{$\>\!\!$}$ or $S_{m}\geq m^{2}\mbox{$\>\!\!$}$, respectively.
@@
2.5 The unconstrained index $g^{*}$
Turning to the verification of the required properties (C1)–(C3) for the $g$-index, we see that (C1) automatically holds due to (7). It is also easy to see that (C3) holds as well. However, the result of Lemma 2 suggests, a bit surprisingly, that property (C2) may fail. For instance, for $\bm{x}=(4)$ we have $g(\bm{x})=1$, but for $\bm{y}=(4,0)$ definition (6) yields $g(\bm{y})=2$.
To salvage (C2), and also to amplify the role of top-cited papers, it was suggested (Egghe, 2006b; Woeginger, 2008b) to lift the constraint $k\leq m$ in the definition of the $g$-index (see (6)) by complementing the citation vector $\bm{x}=(x_{1},\dots,x_{m})$ with additional zeros, as if such fictitious papers have been published but generated no citations: $\bm{x}^{\prime}=(x_{i}^{\prime})=(x_{1},\dots,x_{m},0,\dots)$. We denote this version of $g$ by $g^{*}\mbox{$\>\!\!$}$:
g g $^{*}\mbox{$\>\!\!$}\equiv g^{*}\mbox{$\>\!\!$}(\bm{x})=\max\left\{k\geq 1\colon\sum_{i=1}^{k}x^{\prime}_{i}\geq k^{2}\right\}$ (8)
or, equivalently,
$g^{*}\mbox{$\>\!\!$}=\max\bigl\{k\geq 1\colon S_{k}\geq k^{2}\bigr\},$ (9)
where we define $S_{k}=S_{m}$ for all $k\geq m$.
Comparing definitions (6) and (8), we see that
$g$ $\leq g^{*}\mbox{$\>\!\!$}$,
and moreover, if $g<m$ then g=g $g^{*}\mbox{$\>\!\!$}$. However, the case where $g=m$ may be drastically different.
@@ 2.5 The unconstrained index $g^{*}$
Suppose that $S_{m}\geq(m+1)^{2}\mbox{$\;\!\!$}$. Then $g=m$ but $g^{*}=\displaystyle\bigl\lfloor\sqrt{S_{m}}\bigr\rfloor\geq m+1$.
@@
Note that $g=m$ by Lemma 2. By definition (9) we have
[…]
In turn, this implies
$\sqrt{S_{m}}-1<g^{*}\!\leq\sqrt{S_{m}},$
that is, $g^{*}\!=\displaystyle\bigl\lfloor\sqrt{S_{m}}\bigr\rfloor\geq m+1$, as claimed. ∎
For example, for $\bm{x}=(5,4)$ we have $m=2$, $S_{m}=9$, $g=2$ and $g^{*}\!=3$. A striking real-life example illustrating this situation is the case of John Nash (see Woeginger, 2008b), with the (rounded) citation vector $\bm{x}=(2000,2000,1500,1000,400,250,100,100)$, for which we get $g=8$ but $g^{*}\!=85$.
3 An alternative citation index $\nu$
3.1 Idea and definitions
Trying to reconcile the definition of the $h$-index given by formula (3), with the definition of the $g$-index in (5) by taking into account the actual citations of the top papers, we propose a new citation index called the $\nu$-index, defined as the maximum integer $\nu$ such that the total sum of citation counts of papers with at least $\nu$ citations each is not less than $\nu^{2}\mbox{$\;\!\!$}$. Mathematically, this is expressed as (cf. (3))
$\nu\equiv\nu(\bm{x})=\max\left\{j\geq 1\colon\sum_{i=1}^{m}x_{i}\mbox{$\>\!$}\bm{1}_{\{x_{i}\geq j\}}\geq j^{2}\right\}\!,$ (10)
or, equivalently,
$\nu=\max\bigl\{j\geq 1\colon S_{m_{*\mbox{$\;\!\!$}}(j)}\geq j^{2}\bigr\}.$ (11)
Similarly to (3), the maximum is uniquely defined, noting that the sum in (10) is a decreasing function of $j$, while the right-hand side is strictly increasing. It is also worth pointing out that, unlike $g$ vs. $g^{*}\mbox{$\>\!\!$}$, the $\nu$-index is insensitive to fictitious zeros.
Simple examples show that the value of $\nu$ may be larger than the total number of papers, in contrast with the $h$ and $g$ indexes. For instance, for $\bm{x}=(9,7,1)$ we get $\nu(\bm{x})=4>3$.
@@ 3.1 Idea and definitions
$\bar{\nu}\equiv\bar{\nu}(\bm{x})=\max\left\{1\leq j\leq m\colon\sum_{i=1}^{m}x_{i}\mbox{$\>\!$}\bm{1}_{\{x_{i}\geq j\}}\geq j^{2}\right\}\!,$ (12)
or, equivalently,
$\bar{\nu}=\max\bigl\{1\leq j\leq m\colon S_{m_{*\mbox{$\;\!\!$}}(j)}\geq j^{2}\bigr\}.$